A 20.0L tank contains 0.225 kg of helium at 18 degrees celcius. The molar mass of helium is 4.00 g/mol.

1. How many moles of helium are in the tank?

2. What is the pressure in the tank in Pascals and in atmospheres?


1
Expert's answer
2021-05-12T06:24:11-0400

(a)Given : Volume of tank = 20.0 L

Mass of helium gas = 0.225 Kg = 225

g                         (Since 1 Kg = 1000 g)

And temperature of gas = 18 oC

 

1. The moles of helium in the can is given by,

=> Moles of He = "\frac{Mass of He}{Molar mass of He}=\frac{225}{4.00}"

56.25 mol.



(b) According to ideal gas law,

=> PV = nRT

where P = pressure in atm 

V = volume in L = 20.0 L

n = moles = 56.25 mol.

R = gas constant = 0.0821 atm.L/K.mol

T = temperature in K = T in °C + 273 = 273 + 18 = 291 K


Hence substituting the values we get,

=> P X 20.0 = 56.25 X 0.0821 X 291 

=> P = Pressure in tank = 67.2 atm.

Since 1 atm = 101325 Pa

Hence pressure in tank in Pa = 67.2 X 101325 = 6.81 X 106 Pa approx.